OpenStudy (alina123):

2. Jay borrowed $15,000 for one year from a bank in order to buy a Ducatti motorcycle. He agreed to pay an interest rate of 8% per year. At the end of one year, how much money would Jay pay back to the bank

OpenStudy (ranga):

Interest = P * r * t where P is the principal borrowed = $15,000 r = annual interest rate in decimal = 8/100 = 0.08 t = number of years of the loan = 1 year Interest = 15000 * 0.08 * 1 = ? Add the interest to the principal and that is how much he will owe the bank after one year.

OpenStudy (alina123):

Thank you and i have another question how do you 18 months is equal to 1.5 years

OpenStudy (alina123):

since 1 year is 12 months shouldnt it be 1.6 years for 6 18 minus 12

OpenStudy (ranga):

There are 12 months to a year. 6 months equals 6/12 year = 0.5 year.